OpenStudy (anonymous):

ramone was traveling on the interstate.He noticed the speed signs require that he travel at least 40mph,and at nost 75. Let s stand for Ramone's speed. Write an inequality for this situation.

OpenStudy (anonymous):

Start with each one separately and see if you can combine them.

OpenStudy (anonymous):

\[40 \le x \le 75\]

OpenStudy (anonymous):

Explanation: The part which says "at least 40" means "equal to or more", i.e. \[ \ge \]The part which says "at most 75" means "equal to or less - as long as it is not more", i.e. \[ \le \]So, \[ x \ge 40 \]and \[ x \le 75, \]getting what @gaara438125 said - Note: I just did what @CliffSedge mentioned.

OpenStudy (anonymous):

How's that code-of-conduct lookin' @gaara438125 ?

OpenStudy (anonymous):

yeah, cuz I'm sure she understood an explanation considering there are some things you can't just explain but instead explain, so you find a need to try and look like a "good person" by following the rules, the trying to look smart by failing to even link me the code of conduct, and in the end making yourself look like a fool, and a jerk.

OpenStudy (anonymous):

Know your manners, gaara. And stop pointing fingers at other people before looking at your posts.

OpenStudy (anonymous):

@CliffSedge "Don't be offensive, inappropriate, or creepy! So avoid: Insulting, berating, or otherwise discouraging others."

OpenStudy (anonymous):

and stop harassing me about something completely stupid @Ahaanomegas you telling me to know MY manners when you're focusing me "pointing fingers at me" when he began "pointing fingers" at me....nice hypocrisy

OpenStudy (anonymous):

Nice , 'tu quoque,' @gaara438125 - unfortunately invalid, and unsuccessful.

OpenStudy (anonymous):

could you guys just be quiet?y/n?

OpenStudy (anonymous):

I have a flaw. I am blatantly honest. @sarahlou - I'm with you. I should shut up and this thread should finish. We're making this a boxing ring of words.

OpenStudy (anonymous):

@CliffSedge so, you're allowed to harass me, and not just link me the code of conduct? you can be disrespectful to me, and not vice versa? and i say again, i highly doubt she understood your reasoning on how to solve the question, quite honesty if i didn't know how to do it, you're explanation wouldn't explain anything to me. "Start with each one separately and see if you can combine them." what....tell me WHAT are you combining...?

OpenStudy (anonymous):

^ pitiful.